OpenStudy (aye_my_peeps):

WHAT IS THE SOLUTION TO THIS PROPORTION ? 4/12=10/? A. 3.3 B. 3 C. 30 D. 4.8

rebeccaxhawaii (rebeccaxhawaii):

what is 10/4 then multiply that by12

OpenStudy (isaidavila):

If you multiply both numerator and denominator by 2.5 you'll get the answer

jhonyy9 (jhonyy9):

write 4 10 --- = ---- cross multiplie and calculi the x 12 x

OpenStudy (ineedhelplz):

4/12 = 10/x Answer choice: C)30

jhonyy9 (jhonyy9):

4x=120 x = 120/4 =30 x=30 so you are right sure